Sodium chloride is often used to prevent icy roads and to freeze ice cream. What are the boiling and freezing points of a 0.029m aqueous solution of sodium chloride? (i = 2)

1
Expert's answer
2022-03-25T08:01:05-0400

∆Tb = i × Kb × m = 2 x 0.51°C/m × 0.029 m = 0.03°C

Tsol = 100°C + 0.03°C = 100.03°C

∆Tf = i × Kf × m = 2 × 1.86°C/m × 0.029 m = 0.11°C

Tsol = 0°C - 0.11°C = -0.11°C